OpenStudy (anonymous):

The equation of a line is 2x -2y = 6. What is the slope of the line? Option A: -3 Option B: -1 Option C: 1/3 Option D: 1

OpenStudy (anonymous):

move the function into y = mx + b

OpenStudy (anonymous):

what does that look like?

OpenStudy (anonymous):

I have no idea. I completely forgot how to do that.

OpenStudy (anonymous):

solve for y, do you know how to do that?

OpenStudy (anonymous):

so would you subract y from both sides?

OpenStudy (anonymous):

jk i mean add

OpenStudy (anonymous):

your goal is to get y by itself. so in this equation 2x - 2y = 6 lets move our x away from our y 2x - 2x -2y = 6 - 2x this gives us -2y = 6 - 2x what would be the next step?

OpenStudy (anonymous):

divide by -2y?

OpenStudy (anonymous):

only divide by -2 because we want y to be by itself. so the next step would look like this -2y / -2 = (6 - 2x) / -2 the -2's cancel out on the left. on the right we can fix it to 6/ (-2) - (2x)/(-2) which would reduce down so that y= ??

OpenStudy (anonymous):

y=-3+x?

OpenStudy (anonymous):

correct! so now that you solved for y, you can put the equation into y = mx + b form where m is the slope. y = x - 3 so what would be the slope, m?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Congratulations!! you have found the slope of your line!

OpenStudy (anonymous):

yay thank you!! :))